OpenStudy (trantom):

top 2 answers get a MEDAL :) How would you simplify the inequality 2x-4y < 3/2x+7y

OpenStudy (nikato):

Try to move x to one side of the inequality and y to the other side

OpenStudy (anonymous):

2x-4y<3/2x+7y

OpenStudy (anonymous):

then, subtract 2x from both sides

OpenStudy (trantom):

please type everything in one post... it floods the post please

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

2x-4y < 3/2x+7y Step 1: Move "x" to one side: 2x-4y -3/2x < 3/2x - 3/2 x + 7y 2x-3/2x - 4y < 7y 1/2x - 4y< 7y Step 2: Move "y" to one side 1/2 x < 11y x<22 y is the answer!

OpenStudy (anonymous):

2x-4y < 3/2x+7y +4y +4y 2x<3/2x+11y -3/2x -3/2x 1/2x<11y dived each side by 1/2 x<22y!!

OpenStudy (trantom):

@derk765 i will make a new question and give you the medal on that one k

OpenStudy (anonymous):

It doesn't matter to me :D if you need help I will help, if not I will move on

OpenStudy (trantom):

I made the problem up in about 5 seconds and got the answer in about 3, this question was purely to see who would explain the process well. :)

OpenStudy (anonymous):

lol- im in calculus, have better things to do, so yeah... ttyl!

OpenStudy (trantom):

I'm sort of ishy in calculus xD

OpenStudy (trantom):

as in, I am in(ish) calculus

OpenStudy (anonymous):

meaning?

OpenStudy (trantom):

I'm technically taking algebra but public school is so SLOW I do online stuff. Public school could do all of highschool math in one school year but nope! they make you spend 99% of your time practicing concepts you ALREADY KNOW!!

OpenStudy (anonymous):

ikr- we have worthless teachers

OpenStudy (trantom):

The teachers aren't the problem where i go, its the curiculumm and the few students who have to go slow and slow down everyone else

OpenStudy (trantom):

I mean, you can look at a five minute youtube video and learn the same stuff you take a WEEK to learn in public school.

OpenStudy (anonymous):

that is very true... I take my math online right now

OpenStudy (trantom):

How many years ahead (according to public schools) are you? if you don't want to disclose this because i could subtract and find out your age according to the law of thermodynamics volume II (that last part was gibberish, just random words :)) then by all means restrain your fingers

OpenStudy (anonymous):

well, I'm taking AP calculus as a junior...

OpenStudy (trantom):

three years ahead not bad :)

OpenStudy (trantom):

I'm your biggest fan! (and your only fan... xD)

OpenStudy (trantom):

You're knew here I see

OpenStudy (trantom):

LOL I can't believe I just said that! LOL I meant new omg wow I fail xD

OpenStudy (anonymous):

well, sort of, I only use it when I have algebra questions I don't understand haha

OpenStudy (trantom):

I think thats the first time I have used texting acronyms on this site since I joined..

OpenStudy (anonymous):

:)

OpenStudy (trantom):

your taking algebra II in your school?

OpenStudy (anonymous):

no, I'm taking AP calculus, but I sometimes get stuck on basic algebra because our teachers are that bad

OpenStudy (anonymous):

then I need an outside source

OpenStudy (trantom):

oh, I see

OpenStudy (trantom):

our teachers are good :) our curriculum sucks :( math is still amazing :)

OpenStudy (anonymous):

like tonight i needed help isolating y in y^2+xy+x^2=3

OpenStudy (trantom):

gosh I'm stuck lol

OpenStudy (anonymous):

did you try it?

OpenStudy (trantom):

yea lol I got as far as y^2+xy=3-x^2 then got stuck xD

OpenStudy (anonymous):

you have to complete the square on the LH side

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (trantom):

complete a square? like, y^2=y*y? or what

OpenStudy (anonymous):

no, like this.. http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (trantom):

still not getting it :(

OpenStudy (anonymous):

have you done quadratics?

OpenStudy (trantom):

no lol

OpenStudy (anonymous):

yeah, you have to know how to do those first- it only gets better from there lol

OpenStudy (anonymous):

well, I g2g, so cya later!!!

OpenStudy (trantom):

can you just tell me how to do that problem.... meh. you left wonderful

OpenStudy (anonymous):

:D